Differentiate the following function: f(x)=ex+xe
1
Expert's answer
2020-11-16T19:57:47-0500

"f(x) = e^x + xe^x\\
\textsf{differentiating }e^x\textsf{, we get }e^x"


"\textsf{differentiating } xe^x\textsf{, we use the product rule}\\
\textsf{let u} = x\\
\textsf{let v} = e^x\\
du = 1\\
dv = e^x\\
\textsf{from, }Udv+ Vdu\\
\textsf{we get }xe^x + e^x \textsf{ as the differentiated solution}"

Substituting the differentiated value back into the original question, we have;


"f'(x) = (e^x) + (xe^x + e^x)\\
f'(x) = 2e^x + xe^x\\
\therefore f'(x) = e^x(x+2)"